OpenStudy (anonymous):

Solve the following system of equations using elimination: 4x - 2y = 8 3x - 9y = -54 A. (-3, 5) B. (0, -4) C. (2, 0) D. (6, 8) I got b.

hero (hero):

Hmmm, 3(0) - 9(-4) = -54 36 = -54 False

hero (hero):

I wonder what @wolf1728 is doing. He's been typing for a while now...

OpenStudy (wolf1728):

A) 4x - 2y = 8 B) 3x - 9y = -54 Multiply equation B by -4/3 -4x +12y = +72 Now add this to equation A 4x - 2y = 8 Leaving us with 10y = 80 So now we know y = 8 I think this gives you a good start on solving this.

OpenStudy (anonymous):

4x - 2y = 8 --> (-3) 3x - 9y = -54 -->(4) ------------------- -12x + 6y = - 24 (result of multiplying by -3) 12x - 36y = - 216 (result of multiplying by 4) -------------------add 0 - 30y = - 240 -30y = -240 y = 8 now sub 8 in for y in either of the original equations.. 3x - 9y = -54 3x - 9(8) = -54 3x - 72 = -54 3x = -54 + 72 3x = 18 x = 6 I got d.

OpenStudy (anonymous):

Thanks guys my last answer was D.

OpenStudy (anonymous):

you have to be real careful with elimination problems because it is easy to make mistakes, and if the mistake is made at the beginning of the problem, it messes the whole problem up.

hero (hero):

There are things that can be done to minimize mistakes...such as reducing equations at the beginning. For example. You can divide both sides of the first equation by 2 to get: 2x - y = 4

OpenStudy (anonymous):

so true Hero....